OpenStudy (anonymous):

The line that is the perpendicular bisector of the segment whose endpoints are R(-1, 6) and S(5, 5)

OpenStudy (anonymous):

@austinL

OpenStudy (anonymous):

@ivettef365

OpenStudy (anonymous):

ok i will tell you listen carefully

OpenStudy (anonymous):



OpenStudy (anonymous):

mid point B=(-1+5)/2 and (6+5)/2 which is (2 and 11/2)

OpenStudy (anonymous):

slope of PQ=(5-6)/(5--1)=-1/6 so slope of AB=6

OpenStudy (anonymous):

y-11/2 =6(x-2)